A mass of 3kg hangs 2m above the ground.what is the gravitational potential energy of the mass-earth system at
(1)the ground
(2) a height of 3m above the ground
(3) a height of 4m
1
Expert's answer
2020-05-18T10:30:34-0400

1)


"E=mgh=(3)(10)(0)=0\ J"

2)


"E=mgh=(3)(10)(3)=90\ J"

3)


"E=mgh=(3)(10)(4)=120\ J"